OpenStudy (anonymous):

Question 7 (Essay Worth 5 points) (08.03) Consider the following pair of equations: y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

jhonyy9 (jhonyy9):

do you know the substitution method ?

jhonyy9 (jhonyy9):

y=3x+3 y=x-1 for example from the second equation y=x-1 the value of y substitute in place of y inside first equation - ok? and solve it for x

jhonyy9 (jhonyy9):

so write it please what will get ?

jhonyy9 (jhonyy9):

courage please i will help you

jhonyy9 (jhonyy9):

so than will get x-1=3x+3 can you solve it now for x ?